Kobe completed a certain job in 20 hrs. Together, Kobe and Justin completed the same job in 12 hrs. How long would it take Justin to complete by himself?

Answer by checkley71(8403)

20X/(20+X)=12 CROSS MULTIPLY
20X=12(20+X)
20X=240+12X
20X-12X=240
8X=240
X=240/8
X=30 HRS FOR JUSTIN WORKING ALONE TO COMPLETE THE JOB.
PROOF:
20*30/(20+30)=12
600/50=12
12=12